OpenStudy (anonymous):

Please Help! Medal and Fan A mysterious box is delivered to the dinner party you are attending. The label on the box says that the volume of a box is the function f(x) = x3+3x2-10x (minus) 24. To open the box, you need to identify the correct factors of f(x). Partygoers offer up solutions, and it is your job to find the right ones.

OpenStudy (anonymous):

(x-3)(x+2)(x+4)

OpenStudy (anonymous):

Thank you for the answers but i also need to know how to do it.

OpenStudy (anonymous):

you should try the factors of 24 for x=3 it's true. then you should divide the polinome by (x-3). now you have a quadratic function. x^2+6x+8 can you solve that?

OpenStudy (anonymous):

@marsss Ya

OpenStudy (anonymous):

is it ok? did you understand?

OpenStudy (anonymous):

What do you mean 'for x = 3 its true'

OpenStudy (anonymous):

let x be 3; 3^3+3*3^2-10*3-24=0 i mean x=3 is a solution of this equation

OpenStudy (anonymous):

Okay but how did you get 3

OpenStudy (anonymous):

i tried the factors of 24. for x=1, it's not correct. for x=2, it's not correct. finally for x=3, it's correct lol

OpenStudy (anonymous):

lol okay i understand what you mean, so now you factor out 3 from each monomial now, right?

OpenStudy (anonymous):

yes totally right dude :D

OpenStudy (anonymous):

Okay but you cant factor 3 from 10

OpenStudy (anonymous):

no no no! are you familiar with polinome dividing?

OpenStudy (anonymous):

Kinda

OpenStudy (anonymous):

you should divide f(x) by (x-3)

OpenStudy (anonymous):

so like this \[\frac{ x3 + 3x2 – 10x – 24 }{ x - 3 }\]

OpenStudy (anonymous):

exactly. click the link i've sent. it'll be helpful

OpenStudy (anonymous):

Okay and then once i divide i get \[x^2+6x+8\]

OpenStudy (anonymous):

great! now you know how you solve it

OpenStudy (anonymous):

Okay but what do i do next, because the question is asking for all factors, would i just repeat what we just did except replace the x-3 with something else

OpenStudy (anonymous):

you should solve x^2+6x+8 x^2+6x+8=(x+2)(x+4)

OpenStudy (anonymous):

Okay, so is (x+2) and (x+4) my only factors

OpenStudy (anonymous):

and of course (x-3)

OpenStudy (anonymous):

lol yes, cant forget that one. Well thankyou soooo much!!!

OpenStudy (anonymous):

welcome. good luck